Radio waves from a star, of wavelength 228 m, reac
Answer #1

26degree
Δ=λ2-λ1=λ(1-sinα)

if the first minimum(Δ=λ/2)
occur when θ=26,then
α=180-(θ+90+θ)=90
λ2=Δ/1-sinα
=(228/2)/1-sin90
=319.2
h=λ2sin26

=319.2sin26=139.92
=139.92zin